OpenStudy (gabylovesyou):

The volume of a cylinder is 24 cubic feet. What is the volume of a cone that fits exactly inside the cylinder? Provide an explanation and proof for your answer to receive full credit.

OpenStudy (gabylovesyou):

@jhonyy9 @johnweldon1993

OpenStudy (aum):

Volume of a cylinder = \(\large \pi r^2h\) = 24 cubic feet Volume of a cone = \(\large \frac 13\pi r^2h\) = ?

OpenStudy (gabylovesyou):

8 ?

OpenStudy (aum):

Yes.

OpenStudy (gabylovesyou):

thank you

OpenStudy (aum):

you are welcome.